# Flowr.root – A flow matching based foundation model for joint multi-purpose structure-aware 3D ligand generation and affinity prediction

### v2

@@ Abstract
We present Flowr.root, an -equivariant flow-matching model for pocket-aware 3D ligand generation with joint binding affinity prediction and confidence estimation. The model supports multiple design modes including de novo generation, interaction/pharmacophore-conditional sampling, fragment elaboration, and multi-endpoint affinity prediction (pIC50, pKi, pKd, pEC50). Training combines large-scale ligand libraries with mixed-fidelity protein–ligand complexes, followed by refinement on curated co-crystal datasets and adaptation to project-specific data through parameter-efficient finetuning. Flowr.root achieves state-of-the-art performance in both unconditional 3D molecule and pocket-conditional ligand generation, producing geometrically realistic, low-strain structures with computational efficiency on established benchmark datasets. The integrated affinity prediction module demonstrates superior accuracy on the SPINDR test set and outperforms recent models on the Schrödinger FEP +/OpenFE benchmark while offering substantial speed advantages. As a foundation model, Flowr.root requires continuous parameter-efficient finetuning on project-specific datasets to account for unseen structure-activity landscapes, which we demonstrate yields strong correlation with experimental in-house data. The model’s joint generation and affinity prediction capabilities enable inference-time scaling through importance sampling, effectively steering molecular design toward higher-affinity compounds. Case studies validate this approach: selective CK2 ligand generation against CLK3 shows significant correlation between predicted and quantum-mechanical binding energies, while scaffold elaboration studies on ER and TYK2 demonstrate strong agreement between predicted affinities and QM calculations. By integrating structure-aware generation, affinity estimation, and property-guided sampling within a unified framework, Flowr.root provides a comprehensive foundation for structure-based drug design spanning hit identification through lead optimization.
@@ 1 Introduction
Diffusion- and flow-based generative models have emerged as leading frameworks in machine learning (Ho et al., 2020b; Song et al., 2021b; Rombach et al., 2022). Denoising Diffusion Probabilistic Models (DDPMs) learn a reverse process to transform noise into data samples (Ho et al., 2020b), while score-based models formulate generation via reverse-time stochastic differential equations (SDEs) or their probability-flow ordinary differential equation (ODE) counterparts (Song et al., 2021a). Flow matching, in turn, directly regresses the velocity field along a prescribed probability path to train continuous normalizing flows (Lipman et al., 2023; Liu et al., 2022), with stochastic-interpolant theory offering a unified perspective on diffusion and flow models (Albergo and Vanden-Eijnden, 2023; Albergo et al., 2023). Beyond applications in natural language processing and computer vision, these models are increasingly applied to biology and chemistry, fueling advances in generative chemistry for drug discovery (Hoogeboom et al., 2022; Vignac et al., 2023; Le et al., 2023; Schneuing et al., 2023; Guan et al., 2023; Cremer et al., 2024; Campbell et al., 2024; Irwin et al., 2024; Cremer et al., 2025). In drug discovery, designing small molecules that selectively bind to specific protein targets remains a critical challenge. Diffusion and flow models have proven effective at capturing complex molecular and structural distributions, advancing methods from 3D molecular generation in structure-based drug design (SBDD) to protein–ligand (co-)folding (Abramson et al., 2024; Wohlwend et al., 2025).
@@
One important subfield of AI-driven SBDD is pocket-conditional, structure-aware ligand design. Models like DiffSBDD (Schneuing et al., 2023) and TargetDiff (Guan et al., 2023) have demonstrated that -equivariant diffusion models can generate ligands directly within protein pockets. Follow-up studies have focused on improving chemical validity and pose accuracy (Le et al., 2023; Wang et al., 2024) while coupling pocket conditioning with large-scale pretraining and multi-objective importance sampling to guide generation toward potency and synthesizability (Cremer et al., 2024). Additionally, AI-driven models have increasingly integrated fragment priors into generative pipelines (Imrie et al., 2020; Voloboev, 2024; Zhang et al., 2024; Lee et al., 2025; Guo et al., 2023; Igashov et al., 2024). In fragment-based drug discovery (FBDD), which focuses on hit-to-lead optimization, models typically specialize either in de novo design or fragment-based strategies. However, practical lead optimization requires flexible navigation between these regimes, such as fragment growing or scaffold hopping under interaction constraints, which most current frameworks do not adequately address. Furthermore, many models do not incorporate the geometric constraints imposed by protein pockets. Recent work proposes a unified flow matching approach that integrates de novo, interaction-constrained, and fragment-based generation under pocket conditioning with efficient ODE sampling (Cremer et al., 2025).
@@
However, the effectiveness of generated ligands ultimately depends on their potency, which requires reliable affinity prediction. While experimental validation remains indispensable, computational prioritization during the design phase relies on accurate affinity predictions to effectively navigate target-relevant regions of chemical space. Classical scoring functions, such as AutoDock Vina (Trott and Olson, 2010), Glide (Friesner et al., 2004), and GOLD (Jones et al., 1997), offer computational efficiency but often lack the necessary accuracy for reliable prioritization. Physics-based methods like free energy perturbation (FEP) and absolute binding free energy (ABFE) calculations provide higher precision (Wang et al., 2015; Mey et al., 2020; Mobley and Klimovich, 2012; Alibay et al., 2022; Feng et al., 2022a; Ries et al., 2024), yet their computational cost prohibits application to large-scale generative campaigns, restricting their use to small subsets of candidates where absolute affinity data is critical. Machine learning–based scoring functions improve throughput but suffer from dataset bias and limited generalization (Jiménez et al., 2018). Recent approaches using data augmentation have shown modest improvements (Valsson et al., 2025), though they lack explicit structural information and remain insufficient for robust affinity prediction. The Boltz-2 ligand–protein co-folding model achieves near-FEP accuracy on selected targets with substantial speed advantages (Wohlwend et al., 2025). However, its decoupled architecture—where the affinity head is trained independently from the structural module—creates dependence on structure prediction quality, requires co-folding prior to each affinity prediction, and may limit co-adaptation between protein pocket and ligand geometry, particularly during project-specific fine-tuning.
@@
These limitations motivate a framework that jointly learns structure and affinity while supporting fast, controllable, structure-aware generation. A model that learns the joint probability distribution over ligand geometry, binding pose, and affinity within the protein pocket context would enable efficient, potency-guided ligand generation with on-the-fly ranking capabilities. Crucially, joint training facilitates simple project-specific adaptation through fine-tuning—an essential capability given the distinct constraints of drug discovery campaigns, ranging from ADME/T requirements to R-group and scaffold novelty constraints, diverse assay readouts, and medicinal chemistry heuristics. A fundamental challenge in structure-based generative modeling lies in the inherent disconnect between public-domain training data and project-specific structure-activity relationships (SARs). While models can achieve broad coverage of chemical space through ligand generation, generalizing across novel bioactivity landscapes represents a fundamentally more complex problem. We posit that expecting universal generalization without adaptation is unrealistic; instead, models should function as dynamic companions that continuously refine their understanding of project-specific SARs through sustained interaction with incoming data. This paradigm shift necessitates moving from static models to efficient iterative refinement processes where model utility grows through continuous adaptation. Additionally, techniques such as multi-objective guidance via inference-time importance sampling (Cremer et al., 2024) can help steering generation toward pre-specified, desired properties, while deeper distribution mismatches or activity cliffs can be addressed through direct preference alignment (Schneuing et al., 2025) when suitable data is available.
@@
While large ligand-only resources such as ZINC and PubChem offer abundant chemical diversity (Irwin et al., 2020; Bolton et al., 2011), high-quality protein–ligand complexes with reliable affinity annotations remain scarce and noisy. Databases like PDBBind provide valuable training data (Wang et al., 2005), yet suffer from quality and coverage limitations; curated subsets such as HiQBind address these issues and even increase sample sizes (Wang et al., 2025), yet remain limited in scale. The Plinder dataset (Durairaj et al., 2024) expands scale by broadly curating the PDB, but most entries lack affinity annotations. More recent resources—BindingNet (Zhu et al., 2025), Kinodata-3D (Backenköhler et al., 2024), and especially SAIR (Lemos et al., 2025)—improve chemical space coverage and include affinity data, though they do so with reduced accuracy. This heterogeneous data landscape however, motivates a multi-stage training paradigm: by systematically combining available resources ranked by fidelity, we can exploit their complementary strengths. Large-scale, lower-fidelity data establishes broad chemical space coverage and foundational structural understanding, while subsequent refinement on higher-quality, curated datasets sharpens affinity prediction and structural accuracy. Critically, this approach yields foundation models capable of efficient adaptation to project-specific objectives.
@@
In this work, we present Flowr.root, a foundational framework for structure-based generative chemistry designed to address the continuum from large-scale pre-training through project-specific adaptation. Flowr.root integrates de novo, interaction- and pharmacophore-, as well as fragment-conditioned ligand generation with multi-endpoint affinity prediction and per-sample confidence estimation. Built on the FLOWR.multi backbone (Cremer et al., 2025), the model employs a three-stage training strategy: (1) large-scale pre-training on billions of ligands and millions of protein–ligand complexes with diverse affinity labels, (2) fine-tuning on curated, higher-fidelity datasets, and (3) project-specific adaptation via parameter-efficient fine-tuning and multi-objective guidance through importance sampling. Critically, Flowr.root jointly trains its confidence and affinity prediction modules with structure generation, enabling affinity to actively shape the generative trajectory and facilitating rapid adaptation to project-specific data distributions. This design supports the model’s intended role as a continuously evolving companion throughout early-stage drug discovery campaigns—assisting in motif identification, core expansion, and R-group design through sustained refinement on incoming structure-activity data.
@@ 2 Related Works
Early neural approaches to 3D molecular generation explored autoregressive models that sequentially add coordinates and atom types while maintaining geometric consistency. Symmetry-aware models such as G-SchNet (Gebauer et al., 2020) and its conditional inverse-design follow-up (Gebauer et al., 2022) demonstrated that enforcing symmetries and conditioning signals can substantially improve the validity and controllability of generated 3D structures. Subsequent work proposed explicit autoregressive flows (Luo and Ji, 2022). In parallel, the diffusion modeling paradigm matured from its non-equilibrium thermodynamics roots (Sohl-Dickstein et al., 2015) through score-based (Song et al., 2021a) and variational formulations (Ho et al., 2020a; Kingma et al., 2021). These ideas rapidly translated to molecular geometry: Xu et al. (2022) and Jing et al. (2023) showed that -equivariant denoisers enable high-quality conformation generation by diffusing in Cartesian and torsional spaces, respectively.
@@
The first -equivariant diffusion model for joint continuous coordinate and atom type generation was EDM (Hoogeboom et al., 2022). Follow-ons pushed the design space along several axes: discrete and continuous combination with explicit bond order learning (Vignac et al., 2023), and enhanced denoising learning objectives (Le et al., 2023), establishing strong baselines.
@@
Thereafter, diffusion-based methods pioneered pocket-aware de novo design, also utilizing -equivariant networks. Notably, DiffSBDD (Schneuing et al., 2023) and TargetDiff (Guan et al., 2023) demonstrated that conditional diffusion models can generate diverse, target-specific ligands within the protein pocket, adhering to symmetry constraints and enabling task versatility through sampling controls. More recently, PILOT (Cremer et al., 2024) combined large-scale pretraining, pocket conditioning, and property guidance, highlighting the importance of multi-objective steering (e.g., drug-likeness, synthesizability) under structure constraints. Together, these works established that pocket-aware 3D diffusion can simultaneously respect symmetry, improve pose realism, and support versatile constraints via conditioning and guided sampling.
@@
Meanwhile, flow matching (Lipman et al., 2023) (FM) was proposed refining continuous normalizing flows by directly regressing a time-dependent velocity field that pushes a simple prior to the data distribution, offering faster sampling and flexible priors. SemlaFlow (Irwin et al., 2024) introduced a scalable -equivariant architecture (Semla) trained via flow matching, achieving state-of-the-art unconditional 3D molecule generation with significant speed-ups. Pushing FM into SBDD, FLOWR (Cremer et al., 2025) extended the Semla-style backbones with an dedicated pocket encoder and mixed continuous/categorical FM, supporting multi-mode de novo, interaction-guided, and fragment-based generation in a single model. FLOWR reports large speedups over pocket-diffusion baselines.
@@
Fragment-based drug discovery (FBDD) motivates models that initiate from fragments and perform growth, linking, or merging under pocket constraints. Early deep generative linker design incorporated 3D information into graph models (DeLinker) Imrie et al. (2020), while SyntaLinker and AutoLinker utilized conditional transformers to synthesize linkers directly in the SMILES space, given fragment pairs and constraints Yang et al. (2020b); Feng et al. (2022b). Recent advances have introduced -equivariant models: DiffLinker formulates linker generation as an -equivariant conditional diffusion, explicitly learning 3D geometry between fragment anchors Igashov et al. (2024). For broader medicinal-chemistry workflows, Link-INVENT extends REINVENT with reinforcement learning to optimize linkers for multiple objectives, demonstrated on fragment linking, scaffold hopping, and PROTAC design Guo et al. (2023). STRIFE extracts target-specific pharmacophoric features to steer elaboration in 3D Hadfield et al. (2022), while AutoFragDiff integrates fragment-wise, autoregressive diffusion with pocket conditioning to improve local 3D geometry during growth Ghorbani et al. (2023). For scaffold hopping, DiffHopp employs an -equivariant graph diffusion model tailored for scaffold replacement conditioned on a protein–ligand complex Torge et al. (2023), and TurboHopp accelerates pocket-conditioned 3D scaffold hopping with consistency models and reinforcement learning-based preference optimization Yoo et al. (2024).
@@
Estimating the change in free energy upon binding (, or affinity) accurately remains a cornerstone of structure-enabled small-molecule discovery. Binding affinity is relevant for all early stages of drug discovery, starting from hit identification, where the goal is to find tight and selective binders, through hit-to-lead and lead optimization, where potency must be balanced with absorption, distribution, metabolism, excretion, safety, toxicity, and efficacy considerations. Given the astronomical size of chemical space, computer-aided drug design (CADD) is indispensable to select and prioritize candidates in silico before spending scarce experimental resources (Reymond et al., 2010; Jorgensen, 2004). Classical structure-based approaches to affinity prediction span knowledge-based scoring and physics-based models grounded in molecular mechanics (Liu and Wang, 2015; Gilson and Zhou, 2007). Heuristic docking scores offer speed at the expense of physical rigor. Empirical scoring functions such as AutoDock Vina, Glide, or GOLD remain widely used due to speed, but show inconsistent results between targets Trott and Olson (2010); Friesner et al. (2004); Jones et al. (1997). Semi-empirical and QM / MM scoring have closed part of the gap at an intermediate cost, for example, SQM2.20 achieves DFT-quality affinity estimates in minutes, but only on selected targets (Pecina et al., 2024; Molani et al., 2024). End-point methods such as MM-PBSA and MM-GBSA combine molecular mechanics with continuum solvation to approximate the from MD snapshots at a relatively low cost and remain widely used when throughput is critical (Kollman et al., 2000; Homeyer and Gohlke, 2012; Still et al., 1990; Gohlke and Case, 2004). Alchemical binding free energy methods, absolute (ABFE) and relative (RBFE), trade throughput for accuracy (Aldeghi et al., 2016; Boresch et al., 2003; Gilson et al., 1997; Fu et al., 2022; Feng et al., 2022a; Cournia et al., 2017; Ross et al., 2023a). Modern workflows based on the free energy perturbation theory (FEP) (Zwanzig, 1954) have achieved impressive accuracy on suitable congeneric series (Abel et al., 2017; Wang et al., 2015; Ross et al., 2023b), but remain sensitive to force fields and system preparation.
@@
Machine learning (ML) offers a complementary path to rapid affinity estimation by learning structure–activity relationships directly from data. The early ML scoring functions used interaction fingerprints and hand-made descriptors (Ballester and Mitchell, 2010; Wójcikowski et al., 2019; Kundu et al., 2018; Boyles et al., 2020). Sequence-based CPI / DTA models (e.g., DeepDTA) encode proteins and ligands from 1D inputs to predict binding affinity (Öztürk et al., 2018), while more recent deep architectures, such as 3D convolutional neural networks and graph neural networks, operate more holistically on complex geometry and interaction graphs (Jiménez et al., 2018; Jiang et al., 2021; Karlov et al., 2020; Nguyen et al., 2021; Li et al., 2021; Meli et al., 2022; Valsson et al., 2025).
@@
ML models are typically trained and evaluated on community benchmarks (e.g., CASF) (Li et al., 2018; Su et al., 2019). However, strong in-benchmark performance does not guarantee generalization. Multiple analyses show that models can overfit ligand biases, struggle on out-of-distribution (OOD) targets, or even partially fit to noise (Volkov et al., 2022; Scantlebury et al., 2023; Yang et al., 2020a; Crusius et al., 2025). This limits their reliability in prospective campaigns and underscores the need for approaches that encode biophysical constraints, reduce dataset shortcuts, and validate on OOD benchmarks. Compounding these challenges is data scarcity: structure-based learning ideally requires reliable affinity measurements paired with high-resolution 3D protein–ligand complexes. Although data augmentation is a mainstay in computer vision and NLP (Paulin and Ivašić-Kos, 2023; Pellicer et al., 2023), generating meaningful molecular data that respect stereochemistry, conformational physics, and pocket geometry remains non-trivial. However, combining ChEMBL and PDBBind through comparative complex structure and enhanced template-based modeling resulted in the BindingNet resource, comprising ca. 690k complexes. This significantly densifies the bioactivity landscape compared to the PDBBind alone (Li et al., 2024; Zhu et al., 2025)
@@
A recent advancement is Boltz-2, a co-folding foundation model that predicts complex structures and based on that binding affinity, approaching FEP-level accuracy on certain targets while running orders of magnitude faster, makes large-scale affinity ranking feasible (Wohlwend et al., 2025). Boltz-2’s affinity module couples structural inference with potency prediction, providing a stronger supervisory signal than post hoc scoring and highlighting the value of unified structure–affinity modeling in end-to-end pipelines(Wohlwend et al., 2025).

### v3

@@ Abstract
We present Flowr.root, an -equivariant flow-matching model for pocket-aware 3D ligand generation with joint binding affinity prediction and confidence estimation. The model supports multiple design modes including de novo generation, interaction/pharmacophore-conditional sampling, fragment elaboration, and multi-endpoint affinity prediction (pIC50, pKi, pKd, pEC50). Training combines large-scale ligand libraries with mixed-fidelity protein–ligand complexes, followed by refinement on curated co-crystal datasets and adaptation to project-specific data through parameter-efficient finetuning. Flowr.root achieves state-of-the-art performance in both unconditional 3D molecule and pocket-conditional ligand generation, producing geometrically realistic, low-strain structures with computational efficiency on established benchmark datasets. The integrated affinity prediction module demonstrates superior accuracy on the SPINDR test set and outperforms recent models on the Schrödinger FEP +/OpenFE benchmark while offering substantial speed advantages. As a foundation model, Flowr.root requires continuous parameter-efficient finetuning on project-specific datasets to account for unseen structure-activity landscapes, which we demonstrate yields strong correlation with experimental in-house data. The model’s joint generation and affinity prediction capabilities enable inference-time scaling through importance sampling, effectively steering molecular design toward higher-affinity compounds. Case studies validate this approach: selective CK2 ligand generation against CLK3 shows significant correlation between predicted and quantum-mechanical binding energies, while scaffold elaboration studies on ER, TYK2 and BACE1 demonstrate strong agreement between predicted affinities and QM calculations. By integrating structure-aware generation, affinity estimation, and property-guided sampling within a unified framework, Flowr.root provides a comprehensive foundation for structure-based drug design spanning hit identification through lead optimization.
@@ 1 Introduction
Diffusion- and flow-based generative models have emerged as leading frameworks in machine learning (ho2020ddpm; song2020sde; rombach2022ldm). Denoising Diffusion Probabilistic Models (DDPMs) learn a reverse process to transform noise into data samples (ho2020ddpm), while score-based models formulate generation via reverse-time stochastic differential equations (SDEs) or their probability-flow ordinary differential equation (ODE) counterparts (song2020sde). Flow matching, in turn, directly regresses the velocity field along a prescribed probability path to train continuous normalizing flows (lipman2022flowmatching; liu2022rectified), with stochastic-interpolant theory offering a unified perspective on diffusion and flow models (albergo2022building; albergo2023stochinterp). Beyond applications in natural language processing and computer vision, these models are increasingly applied to biology and chemistry, fueling advances in generative chemistry for drug discovery (model:edm; model:midi; model:eqgatdiff; model:diffsbdd; model:targetdiff; model:pilot; model:multiflow; model:semlaflow; model:flowr). In drug discovery, designing small molecules that selectively bind to specific protein targets remains a critical challenge. Diffusion and flow models have proven effective at capturing complex molecular and structural distributions, advancing methods from 3D molecular generation in structure-based drug design (SBDD) to protein–ligand (co-)folding (folding:alphafold3; boltz2_preprint).
@@
One important subfield of AI-driven SBDD is pocket-conditional, structure-aware ligand design. Models like DiffSBDD (model:diffsbdd) and TargetDiff (model:targetdiff) have demonstrated that -equivariant diffusion models can generate ligands directly within protein pockets. Follow-up studies have focused on improving chemical validity and pose accuracy (model:eqgatdiff; model:molcraft) while coupling pocket conditioning with large-scale pretraining and multi-objective importance sampling to guide generation toward potency and synthesizability (model:pilot). Additionally, AI-driven models have increasingly integrated fragment priors into generative pipelines (imrie2020delinker; voloboev2024fragment2d; zhang2024fraggen; lee2025fragfm; guo2023linkinvent; model:difflinker). In fragment-based drug discovery (FBDD), which focuses on hit-to-lead optimization, models typically specialize either in de novo design or fragment-based strategies. However, practical lead optimization requires flexible navigation between these regimes, such as fragment growing or scaffold hopping under interaction constraints, which most current frameworks do not adequately address. Furthermore, many models do not incorporate the geometric constraints imposed by protein pockets. Recent work proposes a unified flow matching approach that integrates de novo, interaction-constrained, and fragment-based generation under pocket conditioning with efficient ODE sampling (model:flowr).
@@
However, the effectiveness of generated ligands ultimately depends on their potency, which requires reliable affinity prediction. While experimental validation remains indispensable, computational prioritization during the design phase relies on accurate affinity predictions to effectively navigate target-relevant regions of chemical space. Classical scoring functions, such as AutoDock Vina (trott2010vina), Glide (friesner2004glide), and GOLD (jones1997gold), offer computational efficiency but often lack the necessary accuracy for reliable prioritization. Physics-based methods like free energy perturbation (FEP) and absolute binding free energy (ABFE) calculations provide higher precision (wang2015fepplus; mey2020bestpractices; mobley2012alchemical; alibay2022abfe; feng2022abfe; Ries2024_abfe), yet their computational cost prohibits application to large-scale generative campaigns, restricting their use to small subsets of candidates where absolute affinity data is critical. Machine learning–based scoring functions improve throughput but suffer from dataset bias and limited generalization (jimenez2018kdeep). Recent approaches using data augmentation have shown modest improvements (model:aev-plig), though they lack explicit structural information and remain insufficient for robust affinity prediction. The Boltz-2 ligand–protein co-folding model achieves near-FEP accuracy on selected targets with substantial speed advantages (boltz2_preprint). However, its decoupled architecture—where the affinity head is trained independently from the structural module—creates dependence on structure prediction quality, requires co-folding prior to each affinity prediction, and may limit co-adaptation between protein pocket and ligand geometry, particularly during project-specific fine-tuning.
@@
These limitations motivate a framework that jointly learns structure and affinity while supporting fast, controllable, structure-aware generation. A model that learns the joint probability distribution over ligand geometry, binding pose, and affinity within the protein pocket context would enable efficient, potency-guided ligand generation with on-the-fly ranking capabilities. Crucially, joint training facilitates simple project-specific adaptation through fine-tuning—an essential capability given the distinct constraints of drug discovery campaigns, ranging from ADME/T requirements to R-group and scaffold novelty constraints, diverse assay readouts, and medicinal chemistry heuristics. A fundamental challenge in structure-based generative modeling lies in the inherent disconnect between public-domain training data and project-specific structure-activity relationships (SARs). While models can achieve broad coverage of chemical space through ligand generation, generalizing across novel bioactivity landscapes represents a fundamentally more complex problem. We posit that expecting universal generalization without adaptation is unrealistic; instead, models should function as dynamic companions that continuously refine their understanding of project-specific SARs through sustained interaction with incoming data. This paradigm shift necessitates moving from static models to efficient iterative refinement processes where model utility grows through continuous adaptation. Additionally, techniques such as multi-objective guidance via inference-time importance sampling (model:pilot) can help steering generation toward pre-specified, desired properties, while deeper distribution mismatches or activity cliffs can be addressed through direct preference alignment (drugflow) when suitable data is available.
@@
While large ligand-only resources such as ZINC and PubChem offer abundant chemical diversity (dataset:zinc; dataset:pubchem), high-quality protein–ligand complexes with reliable affinity annotations remain scarce and noisy. Databases like PDBBind provide valuable training data (dataset:pdbbind), yet suffer from quality and coverage limitations; curated subsets such as HiQBind address these issues and even increase sample sizes (dataset:hiqbind), yet remain limited in scale. The Plinder dataset (dataset:plinder) expands scale by broadly curating the PDB, but most entries lack affinity annotations. More recent resources—BindingNet (dataset:bindingnetv2), Kinodata-3D (dataset:kinodata), and especially SAIR (dataset:sair)—improve chemical space coverage and include affinity data, though they do so with reduced accuracy. This heterogeneous data landscape however, motivates a multi-stage training paradigm: by systematically combining available resources ranked by fidelity, we can exploit their complementary strengths. Large-scale, lower-fidelity data establishes broad chemical space coverage and foundational structural understanding, while subsequent refinement on higher-quality, curated datasets sharpens affinity prediction and structural accuracy. Critically, this approach yields foundation models capable of efficient adaptation to project-specific objectives.
@@
In this work, we present Flowr.root, a foundational framework for structure-based generative chemistry designed to address the continuum from large-scale pre-training through project-specific adaptation. Flowr.root integrates de novo, interaction- and pharmacophore-, as well as fragment-conditioned ligand generation with multi-endpoint affinity prediction and per-sample confidence estimation. Built on the FLOWR.multi backbone (model:flowr), the model employs a three-stage training strategy: (1) large-scale pre-training on billions of ligands and millions of protein–ligand complexes with diverse affinity labels, (2) fine-tuning on curated, higher-fidelity datasets, and (3) project-specific adaptation via parameter-efficient fine-tuning and multi-objective guidance through importance sampling. Critically, Flowr.root jointly trains its confidence and affinity prediction modules with structure generation, enabling affinity to actively shape the generative trajectory and facilitating rapid adaptation to project-specific data distributions. This design supports the model’s intended role as a continuously evolving companion throughout early-stage drug discovery campaigns—assisting in motif identification, core expansion, and R-group design through sustained refinement on incoming structure-activity data.
@@ 2 Related Works
Early neural approaches to 3D molecular generation explored autoregressive models that sequentially add coordinates and atom types while maintaining geometric consistency. Symmetry-aware models such as G-SchNet (model:gschnet) and its conditional inverse-design follow-up (model:cgschnet) demonstrated that enforcing symmetries and conditioning signals can substantially improve the validity and controllability of generated 3D structures. Subsequent work proposed explicit autoregressive flows (model:autoregressive_flow). In parallel, the diffusion modeling paradigm matured from its non-equilibrium thermodynamics roots (diffusion:original) through score-based (song2020sde) and variational formulations (diffusion:ddpm; kingma2021vdm). These ideas rapidly translated to molecular geometry: xu2022geodiff and model:torsional_diffusion showed that -equivariant denoisers enable high-quality conformation generation by diffusing in Cartesian and torsional spaces, respectively.
@@
The first -equivariant diffusion model for joint continuous coordinate and atom type generation was EDM (model:edm). Follow-ons pushed the design space along several axes: discrete and continuous combination with explicit bond order learning (model:midi), and enhanced denoising learning objectives (model:eqgatdiff), establishing strong baselines.
@@
Thereafter, diffusion-based methods pioneered pocket-aware de novo design, also utilizing -equivariant networks. Notably, DiffSBDD (model:diffsbdd) and TargetDiff (model:targetdiff) demonstrated that conditional diffusion models can generate diverse, target-specific ligands within the protein pocket, adhering to symmetry constraints and enabling task versatility through sampling controls. More recently, PILOT (model:pilot) combined large-scale pretraining, pocket conditioning, and property guidance, highlighting the importance of multi-objective steering (e.g., drug-likeness, synthesizability) under structure constraints. Together, these works established that pocket-aware 3D diffusion can simultaneously respect symmetry, improve pose realism, and support versatile constraints via conditioning and guided sampling.
@@
Meanwhile, flow matching (lipman2022flowmatching) (FM) was proposed refining continuous normalizing flows by directly regressing a time-dependent velocity field that pushes a simple prior to the data distribution, offering faster sampling and flexible priors. SemlaFlow (model:semlaflow) introduced a scalable -equivariant architecture (Semla) trained via flow matching, achieving state-of-the-art unconditional 3D molecule generation with significant speed-ups. Pushing FM into SBDD, FLOWR (model:flowr) extended the Semla-style backbones with an dedicated pocket encoder and mixed continuous/categorical FM, supporting multi-mode de novo, interaction-guided, and fragment-based generation in a single model. FLOWR reports large speedups over pocket-diffusion baselines.
@@
Fragment-based drug discovery (FBDD) motivates models that initiate from fragments and perform growth, linking, or merging under pocket constraints. Early deep generative linker design incorporated 3D information into graph models (DeLinker) imrie2020delinker, while SyntaLinker and AutoLinker utilized conditional transformers to synthesize linkers directly in the SMILES space, given fragment pairs and constraints yang2020syntalink; feng2022syntalinkhybrid. Recent advances have introduced -equivariant models: DiffLinker formulates linker generation as an -equivariant conditional diffusion, explicitly learning 3D geometry between fragment anchors model:difflinker. For broader medicinal-chemistry workflows, Link-INVENT extends REINVENT with reinforcement learning to optimize linkers for multiple objectives, demonstrated on fragment linking, scaffold hopping, and PROTAC design guo2023linkinvent. STRIFE extracts target-specific pharmacophoric features to steer elaboration in 3D hadfield2022strife, while AutoFragDiff integrates fragment-wise, autoregressive diffusion with pocket conditioning to improve local 3D geometry during growth ghorbani2023autofragdiff. For scaffold hopping, DiffHopp employs an -equivariant graph diffusion model tailored for scaffold replacement conditioned on a protein–ligand complex torge2023diffhopp, and TurboHopp accelerates pocket-conditioned 3D scaffold hopping with consistency models and reinforcement learning-based preference optimization yoo2024turbohopp.
@@
Estimating the change in free energy upon binding (, or affinity) accurately remains a cornerstone of structure-enabled small-molecule discovery. Binding affinity is relevant for all early stages of drug discovery, starting from hit identification, where the goal is to find tight and selective binders, through hit-to-lead and lead optimization, where potency must be balanced with absorption, distribution, metabolism, excretion, safety, toxicity, and efficacy considerations. Given the astronomical size of chemical space, computer-aided drug design (CADD) is indispensable to select and prioritize candidates in silico before spending scarce experimental resources (reymond2010chemicalspace; jorgensen2004manyroles). Classical structure-based approaches to affinity prediction span knowledge-based scoring and physics-based models grounded in molecular mechanics (liu2015classification; gilson2007calculation). Heuristic docking scores offer speed at the expense of physical rigor. Empirical scoring functions such as AutoDock Vina, Glide, or GOLD remain widely used due to speed, but show inconsistent results between targets trott2010vina; friesner2004glide; jones1997gold. Semi-empirical and QM / MM scoring have closed part of the gap at an intermediate cost, for example, SQM2.20 achieves DFT-quality affinity estimates in minutes, but only on selected targets (pecina2024sqm220; molani2024qmmm). End-point methods such as MM-PBSA and MM-GBSA combine molecular mechanics with continuum solvation to approximate the from MD snapshots at a relatively low cost and remain widely used when throughput is critical (kollman2000accresearch; homeyer2012mmpbsa; still1990gb; gohlke2004rasraf). Alchemical binding free energy methods, absolute (ABFE) and relative (RBFE), trade throughput for accuracy (aldeghi2016_abfe; boresch2003abfe; gilson1997statmech; Fu2022_abfe; feng2022abfe; cournia2017rbfe; Ross2023_rbfe). Modern workflows based on the free energy perturbation theory (FEP) (zwanzig1954perturbation) have achieved impressive accuracy on suitable congeneric series (abel2017enhanced; wang2015fepplus; ross2023maximal), but remain sensitive to force fields and system preparation.
@@
Machine learning (ML) offers a complementary path to rapid affinity estimation by learning structure–activity relationships directly from data. The early ML scoring functions used interaction fingerprints and hand-made descriptors (ballester2010rfscore; wojcikowski2019plec; kundu2018fundamental; boyles2020learningligand). Sequence-based CPI / DTA models (e.g., DeepDTA) encode proteins and ligands from 1D inputs to predict binding affinity (ozturk2018deepdta), while more recent deep architectures, such as 3D convolutional neural networks and graph neural networks, operate more holistically on complex geometry and interaction graphs (jimenez2018kdeep; jiang2021interactiongraphnet; karlov2020graphdelta; nguyen2021graphdta; li2021sign; meli2022deepreview; model:aev-plig).
@@
ML models are typically trained and evaluated on community benchmarks (e.g., CASF) (li2018casf2013; su2019casf2016). However, strong in-benchmark performance does not guarantee generalization. Multiple analyses show that models can overfit ligand biases, struggle on out-of-distribution (OOD) targets, or even partially fit to noise (volkov2022frustration; scantlebury2023generalizability; yang2020predictingorpretending; crusius2025fittingnoise). This limits their reliability in prospective campaigns and underscores the need for approaches that encode biophysical constraints, reduce dataset shortcuts, and validate on OOD benchmarks. Compounding these challenges is data scarcity: structure-based learning ideally requires reliable affinity measurements paired with high-resolution 3D protein–ligand complexes. Although data augmentation is a mainstay in computer vision and NLP (paulin2023syntheticcv; pellicer2023nlpda), generating meaningful molecular data that respect stereochemistry, conformational physics, and pocket geometry remains non-trivial. However, combining ChEMBL and PDBBind through comparative complex structure and enhanced template-based modeling resulted in the BindingNet resource, comprising ca. 690k complexes. This significantly densifies the bioactivity landscape compared to the PDBBind alone (dataset:bindingnetv1; dataset:bindingnetv2)
@@
A recent advancement is Boltz-2, a co-folding foundation model that predicts complex structures and based on that binding affinity, approaching FEP-level accuracy on certain targets while running orders of magnitude faster, makes large-scale affinity ranking feasible (boltz2_preprint). Boltz-2’s affinity module couples structural inference with potency prediction, providing a stronger supervisory signal than post hoc scoring and highlighting the value of unified structure–affinity modeling in end-to-end pipelines(boltz2_preprint).